im doing a scale project for school. i dont knw how to put my measurements in proportions for length, width, and depth. could you plz give me examples like make up measurements for length, width, depth and put them in proportions. plz. and thankyou for helping.
Answer by checkley77(12844)

Given you have a barn that is 28ft by 16ft by 12ft. (length width & height).
You want a scale of 1 in. for 4 ft.
28/4=7 inches
16/4=4 inches
12/4=3 inches
So the barn size of 26:16:12 ft.
Becomes 7:4:3 inches.